OpenStudy (anonymous):

What is the largest number that can be obtained as the product of two or more positive integers that add up to 20?

hero (hero):

Should be 10*10 = 100 10 + 10 = 20

hero (hero):

The trick is to always divide in half then use the closest integers

OpenStudy (anonymous):

3*3*3*3*3*3*2=1458 3+3+3+3+3+3+2=20

hero (hero):

Oh, well, if that's what they meant then okay.

OpenStudy (anonymous):

it says two or more positive integers so i assume thats fine

hero (hero):

I didn't see the "or more" part

OpenStudy (anonymous):

Maybe.. 2x2x2x2x2x2x2x2x2x2= 1048576?

OpenStudy (anonymous):

This has one integer, but is a largest answer.. can someone please confirm?

hero (hero):

2^10 = 1024 I don't know how you arrived at your result.

hero (hero):

coug18's solution should be correct.